OpenStudy (anonymous):

x^2 - 4x- 8 =0 How does it work?

OpenStudy (anonymous):

x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} for ax^2 +bx +c

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

thanks for the answers =))))))